OpenStudy (anonymous):

given that X ~ N(m,5^2) and that P(X>63) = 0.367 find m ? you are aloud to use a calculator but i dont know how to find m, without guess and check using a casio class pad calc. please help

OpenStudy (dumbcow):

i will assume this is using the normal distribution where m is the mean and 25 is stnd dev ?

OpenStudy (anonymous):

yep you are correct exept 5 is the std deviation and 5^2 is the variance

OpenStudy (dumbcow):

oh ok...well you need to look up the Z-score also note that most tables only give probabilities for "less than" so change it like this P(X>63) = 1-P(X<=63) so look up 1-.367 then use z-score formula \[Z = \frac{X-mean}{std dev}\]

OpenStudy (anonymous):

hmm what do you mean by 'most tables'? and what would the X be? 1-0.367 ?

OpenStudy (dumbcow):

most stat textbooks provide standard normal probability tables to look up probabilities and z-scores... http://en.wikipedia.org/wiki/Standard_normal_table#Cumulative_table

OpenStudy (anonymous):

oh wow

OpenStudy (dumbcow):

X = 63 the probability of being less than 63 is (1-0.367)

OpenStudy (anonymous):

ohhh okay! i think i got ya thanks!

OpenStudy (dumbcow):

no problem :)

OpenStudy (dumbcow):

i get m = 61.3

OpenStudy (anonymous):

thats correct :)